OpenStudy (anonymous):

Will Medal! cos(a/2) = +/- sqrt (1+cos(a)/2) for all values of a. A. True B. False

OpenStudy (anonymous):

\[\cos(\frac{ a }{ 2}) = \pm \sqrt{\frac{ 1+cosa }{ 2 }}\]

OpenStudy (anonymous):

I think the answer might be true.

OpenStudy (anonymous):

Yeah it is true.